# Detecting and Evaluating Medical Hallucinations in Large Vision Language Models

### v1

@@
Abstract
Large Vision Language Models (LVLMs) are increasingly integral to healthcare applications, including medical visual question answering and imaging report generation. While these models inherit the robust capabilities of foundational Large Language Models (LLMs), they also inherit susceptibility to hallucinations—a significant concern in high-stakes medical contexts where the margin for error is minimal. However, currently, there are no dedicated methods or benchmarks for hallucination detection and evaluation in the medical field. To bridge this gap, we introduce Med-HallMark, the first benchmark specifically designed for hallucination detection and evaluation within the medical multimodal domain. This benchmark provides multi-tasking hallucination support, multifaceted hallucination data, and hierarchical hallucination categorization. Furthermore, we propose the MediHall Score, a new medical evaluative metric designed to assess LVLMs’ hallucinations through a hierarchical scoring system that considers the severity and type of hallucination, thereby enabling a granular assessment of potential clinical impacts. We also present MediHallDetector, a novel Medical LVLM engineered for precise hallucination detection, which employs multitask training for hallucination detection. Through extensive experimental evaluations, we establish baselines for popular LVLMs using our benchmark. The findings indicate that MediHall Score provides a more nuanced understanding of hallucination impacts compared to traditional metrics and demonstrate the enhanced performance of MediHallDetector. We hope this work can significantly improve the reliability of LVLMs in medical applications. All resources of this work will be released soon.
@@ 1 Introduction
To address these issues and enable researchers to evaluate LVLMs’ medical outputs reasonably, we propose solutions from three dimensions: data, evaluation metrics and detection methods. Firstly, we introduce a hierarchical categorization of hallucinations specific to the medical domain and develop Med-HallMark, the first benchmark for hallucination detection in medical multimodal fields which provides multi-tasking hallucination support, multifaceted hallucination data, and hierarchical hallucination categorization. Furthermore, we propose the MediHall Score, a new evaluation metric specifically designed for the medical domain, which calculates the hallucination score of the LVLM outputs through hierarchical categorization, providing an intuitive numerical representation of the rationality of the medical texts. Finally, we present MediHallDetector, the first multimodal medical hallucination detection model designed to detect hallucinations in model output texts with fine granularity. It employs unique design features and customized training methods to enhance its scalability and hallucination detection capabilities. We not only provide the baseline performance of the most popular LVLMs with medical capacity on the Med-HallMark but also demonstrate the rationality of the medical multimodal hallucination detection method in this paper as well as the superiority of the hallucination detection model MediHallDetector.
In general, the main contributions of this paper are as follows: (i) We introduce the first benchmark dedicated to hallucination detection in the medical domain, Med-HallMark, and provide baselines for various LVLMs. (ii) We propose the first hallucination detection model, MediHallDetector, and demonstrate its superiority through extensive experiments. (iii) We present a new hallucination evaluation metric, MediHall Score, and show its effectiveness relative to traditional metrics through qualitative and quantitative analysis.
@@ 3.2 Multifaceted Hallucination Data
Our benchmark includes multifaceted hallucination data, detailed across three primary dimensions: ground truth (GT), LVLM outputs for prompts and annotations of the LVLM-generated content. The GT provide a reliable standard against which to evaluate LVLM performance. The LVLM outputs for prompts may be correct or may contain hallucinations. Med-HallMark includes fine-grained annotations for all LVLM responses, detailing both the type of hallucination and its correctness. In fine-grained single-dimension VQA scenarios, each response is labeled with a single hallucination category. In coarse-grained multi-dimension IRG scenarios, the LVLM outputs are segmented into sentences and annotated at the sentence level. This detailed annotation process allows for a thorough evaluation of model performance across different medical tasks.
@@ 3.3 Hierarchical Hallucination Categorization
To address this gap, we propose a novel hierarchical method specifically designed for medical text hallucinations, which classifies hallucinations based on the severity of their impact on clinical diagnosis or decision-making, as illustrated in Figure 1. We finely categorize the sentence-level hallucination outputs of LVLMs into five distinct levels:
@@
Attribute Hallucinations: These manifest as incorrect judgments or descriptions of the size, shape, location, and number of organs and pathologies and affect diagnostic accuracy to some extent.
Prompt-induced Hallucinations: These hallucinations are induced by prompts containing confused information, often arising from a lack of plausibility or factuality in the prompt and testing the model’s robustness in specific contexts.
@@ 3.4 Construction of the Benchmark
For images from the VQA dataset, we designed a series of questions based on six dimensions: modality, plane, shape, size, organ, location, and pathology, following the methods proposed in [3]. These original questions are considered conventional medical questions $Q_{Conv}$. We then used the (SOTA) model, LLaVA-Med-pretrained [11], to infer answers to $Q_{Conv}$. The generated answers $A_{Conv}$ were manually evaluated. Correct answers $A_{Conv}^{true}$ were directly used as ground truth, while incorrect answers $A_{Conv}^{false}$ were treated as hallucinatory outputs and their ground truths were re-annotated to expedite the annotation process. Incorrect questions $Q_{Conv}^{false}$ were annotated with the hallucination categories introduced in subsection 3.3. To quickly expand the dataset, we utilized the GPT-3.5 API to rewrite $Q_{Conv}$ questions, ensuring the questions’ form did not alter the ground truth, resulting in $Q_{Conv}^{true^{\prime}}$ and $Q_{Conv}^{false^{\prime}}$.
@@
In the IRG scenario, we sampled 1800 images and their corresponding medical reports from the MIMIC-test and OpenI datasets (sampling methods are detailed in the Supplementary material). To generate concise, de-identified medical reports containing complete key information as the dataset’s ground truth, we processed the medical reports by removing sentences comparing the patient’s previous medical history to ensure the accuracy of independent data. We also de-identified the reports by removing patient-specific, doctor-specific, and visit-specific information. After cleaning the data, we concatenated the Findings and Impressions sections of the reports to form the ground truth for the medical image report generation task.
@@ 5 MediHallDetector
Figure 2: Visualization of MediHalldetector related information. (a) Model structure, SFT process and inference objective of MediHalldetector. (b) Examples of questions, LVLM answers and $GT$ for different types of tasks. (c) Comparison of three rounds of evaluation agreement and average inference time for different evaluation models. (d) Comparison of different evaluation models’ agreement with human evaluation preferences in different hallucination texts.
@@
Framework of MediHallDetector: MediHallDetector is built upon the fundamental architecture of the LLaVA model, incorporating a dual-layer fully connected network with GELU activation functions as the connector. The initial weights for MediHallDetector were taken from LLaVA1.5-7B[16] to leverage its robust foundational capabilities.
Data Preparation: The training data for MediHallDetector consists of three main categories: traditional medical image-text task data, Med-HallMark data, and specific hallucination evaluation instruction pair data. Traditional medical image-text task data, sourced from SLAKE, VQA-RAD, MIMIC-Test and OpenI datasets shown in Figure 1, helps adapt the general LVLM to the medical domain. Med-HallMark data improves the model’s robustness and prevents the assessment model from being induced to produce hallucinations by the hallucination output itself due to the problem of not having seen the source domain. The hallucination evaluation data, manually annotated medical text described in Section 3, helps the model understand different types of hallucinations and follow the instructions.
@@ 6 Experiment and Discussion
Model Med-VQA Task IRG Task
BS MR R-1 R-2 R-L BLEU
@@
ACC BS MR R-1 R-2 R-L BLEU
@@
BLIP2 47.97 16.15 18.98 6.03 17.13 3.46 0.52 0.27 33.05 4.88 10.17 1.49 7.66 0.15
InstructBLIP-7b 35.99 7.47 6.08 0.59 5.30 1.03 0.58 0.34 47.49 13.98 17.56 2.31 13.60 0.73 0.84
InstructBLIP-13b 36.02 7.59 6.13 0.58 5.32 1.00 0.58 0.34 47.47 13.93 17.54 2.34 13.61 0.74 0.84
LLaVA1.5-7b 54.89 28.33 23.52 9.30 21.16 4.60 0.54 0.28 47.93 11.24 18.77 2.78 14.64 0.72 0.79
LLaVA1.5-13b 52.82 25.98 21.52 8.20 19.38 4.18 0.51 0.23 47.96 11.80 18.35 2.40 14.49 0.67 0.81
LLaVA-Med(SF) 36.67 8.80 91.17 1.40 9.10 0.03 0.61 0.39 30.49 0.61 0.27 0.00 0.27 0.14
LLaVA-Med(RF) 35.25 6.91 6.34 1.49 6.09 0.54 0.57 0.32 33.28 2.33 2.73 0.32 2.45 0.02
LLaVA-Med(PF) 33.32 3.27 2.85 0.58 2.68 0.06 0.59 0.35 37.67 1.25 2.24 0.05 2.06 0.09
mPLUG-Owl2 55.11 29.39 22.25 8.38 19.77 3.43 0.50 0.24 64.49 40.11 32.00 13.84 28.50 6.79 0.81
XrayGPT 44.40 14.00 10.66 1.17 9.89 0.37 0.36 0.02 62.62 25.96 27.94 6.59 22.15 3.26 0.79
MiniGPT4 42.93 12.93 11.14 1.60 10.19 0.39 0.38 0.09 46.43 10.27 15.37 1.75 12.63 0.53 0.88
RadFM 43.84 11.81 11.31 2.16 10.68 1.55 0.56 0.32 41.18 3.41 5.88 0.76 4.79 0.05
@@ 6.1 Baseline Results on the Med-HallMark
On the IGR task, mPLUG-Owl2 outperforms the other baseline models on most conventional metrics. LLaVA-Meds yields the worst result. The intuitive examples are reflected in the ROUGE-1/2/L with scores close to 0%. A plausible explanation is that LLaVA-Med is full-parameters fine-tuned on the Q&A dataset with short answers, and its inability to perform the IRG task, which requires the giving of long contextual responses. BertScore is considered to be in favourable agreement with human judgment compared to other metrics. In this case, the scores of 64.49% and 62.62% for mPLUG-Owl2 and XrayGPT, respectively. In contrast, BLIP and LLaVA1.5 families achieve comparable performance due to the BertScore of around 47%.
On the MediHall Score metric for assessing hallucinations, we perform corresponding analyses based on different tasks. Specifically, the LLaVA-Med series achieves higher results on the Med-VQA task, which arrive at 0.61, 0.57, and 0.69, respectively, suggesting the production of more correct content and fewer hallucinatory states.
On the IRG task, the LLaVA-Med series, BLIP2, and RadFM cannot produce a computable MediHall Score since the generation format of these models is not suitable for reporting generation scenarios with contextual reasoning properties. Furthermore, MiniGPT4 exhibits favorable performance with the highest score of 0.88.
Figure 3: Comparison of different Models on hallucination types.
Table 2: Comparison of different models in Med-HallMark on $Q_{Conv}^{true^{\prime}}$ ($\clubsuit$’) and $Q_{Inconfi}$ ($\spadesuit$’).
@@ 6.2 Comparison between different Metrics
Table 3 quantitatively illustrates the strengths and weaknesses of different metrics when evaluating LVLM responses. In various question types, LVLM responses may be correct or exhibit hallucinations. As shown in Figure 2(b), regardless of whether the LVLM’s answer $A$ aligns with the ground truth $GT$, the METEOR metric shown in Table 3 fails to directly reflect this alignment.
Table 3: Comparison between traditional Metrics and MediHall Score. Each metric corresponds to QA pairs with the same ID in Figure 2(b)
@@
ROUGE score evaluates the presence of matching n-grams or subsequences between two texts. ROUGE-1/2/L matches single words, bigrams, and the longest common subsequence, respectively. Although this can detect some content alignment between $A$ and $GT$, it is prone to extreme cases. For example, in the confidence-weakening question shown in Figure 2(b), the model correctly identifies the largest organ in the image as the lung. However, the $A$Lung." fails to match "lung" in $GT$ due to the punctuation, resulting in a failed direct match. Additionally, because $A$ contains only one word, ROUGE-2/L cannot be computed. BLEU has similar issues; without any shared n-grams or subsequences between $A$ and GT, BLEU also scores zero. While BLEU does account for significant length differences between A and GT, making it somewhat more versatile than ROUGE, it still weakly measures factual correctness.
6.3 Evaluations of the MediHallDetector’s Performance

### v2

@@
Detecting and Evaluating Medical Hallucinations in Large Vision Language Models
Jiawei Chen1,2,3222Equal first contributions. §Equal second contributions. Corresponding authors.   Dingkang Yang1,3222Equal first contributions. §Equal second contributions. Corresponding authors.
Tong Wu244footnotemark: 4Yue Jiang1,344footnotemark: 4Xiaolu Hou1,344footnotemark: 4
Mingcheng Li1,3Shunli Wang1,2,3Dongling Xiao2
Ke Li2Lihua Zhang1,311footnotemark: 1
1Academy for Engineering and Technology, Fudan University
2Tencent Youtu Lab
3Cognition and Intelligent Technology Laboratory
Large Vision Language Models (LVLMs) are increasingly integral to healthcare applications, including medical visual question answering and imaging report generation. While these models inherit the robust capabilities of foundational Large Language Models (LLMs), they also inherit susceptibility to hallucinations—a significant concern in high-stakes medical contexts where the margin for error is minimal. However, currently, there are no dedicated methods or benchmarks for hallucination detection and evaluation in the medical field. To bridge this gap, we introduce Med-HallMark, the first benchmark specifically designed for hallucination detection and evaluation within the medical multimodal domain. This benchmark provides multi-tasking hallucination support, multifaceted hallucination data, and hierarchical hallucination categorization. Furthermore, we propose the MediHall Score, a new medical evaluative metric designed to assess LVLMs’ hallucinations through a hierarchical scoring system that considers the severity and type of hallucination, thereby enabling a granular assessment of potential clinical impacts. We also present MediHallDetector, a novel Medical LVLM engineered for precise hallucination detection, which employs multitask training for hallucination detection. Through extensive experimental evaluations, we establish baselines for popular LVLMs using our benchmark. The findings indicate that MediHall Score provides a more nuanced understanding of hallucination impacts compared to traditional metrics and demonstrate the enhanced performance of MediHallDetector. We hope this work can significantly improve the reliability of LVLMs in medical applications.  Last Update Date: December 10, 2024 Correspondence: chenjiawei22@m.fudan.edu.cn Code: https://github.com/ydk122024/Med-HallMark
@@ 1  Introduction
To address these issues and enable researchers to evaluate LVLMs’ medical outputs reasonably, we propose solutions from three dimensions: data, evaluation metrics and detection methods. Firstly, we introduce a hierarchical categorization of hallucinations specific to the medical domain and develop Med-HallMark, the first benchmark for hallucination detection in medical multimodal fields, which provides multi-tasking hallucination support, multifaceted hallucination data, and hierarchical hallucination categorization. Furthermore, we propose the MediHall Score, a new evaluation metric specifically designed for the medical domain, which calculates the hallucination score of the LVLM outputs through hierarchical categorization, providing an intuitive numerical representation of the rationality of the medical texts. Finally, we present MediHallDetector, the first multimodal medical hallucination detection model designed to detect hallucinations in model output texts with fine granularity. It employs unique design features and customized training methods to enhance its scalability and hallucination detection capabilities. We not only provide the baseline performance of the most popular LVLMs with medical capacity on the Med-HallMark but also demonstrate the rationality of the medical multimodal hallucination detection method in this paper as well as the superiority of the hallucination detection model MediHallDetector.
In general, the main contributions of this paper are as follows:
(i) We introduce the first benchmark dedicated to hallucination detection in the medical domain, Med-HallMark, and provide baselines for various LVLMs.
(ii) We propose the first hallucination detection model, MediHallDetector, and demonstrate its superiority through extensive experiments.
(iii) We present a new hallucination evaluation metric, MediHall Score, and show its effectiveness relative to traditional metrics through qualitative and quantitative analysis.
@@ 3.2  Multifaceted Hallucination Data
Our benchmark includes multifaceted hallucination data, detailed across three primary dimensions: ground truth (GT), LVLM outputs for prompts, and annotations of the LVLM-generated content. The GT provide a reliable standard against which to evaluate LVLM performance. The LVLM outputs for prompts may be correct or may contain hallucinations. Med-HallMark includes fine-grained annotations for all LVLM responses, detailing both the type of hallucination and its correctness. In fine-grained single-dimension VQA scenarios, each response is labeled with a single hallucination category. In coarse-grained multi-dimensional IRG scenarios, the LVLM outputs are segmented into sentences and annotated at the sentence level. This detailed annotation process allows for a thorough evaluation of model performance across different medical tasks.
@@ 3.3  Hierarchical Hallucination Categorization
To address this gap, we propose a novel hierarchical method specifically designed for medical text hallucinations, which classifies hallucinations based on the severity of their impact on clinical diagnosis or decision-making, as illustrated in Figure 1. With the assistance of experienced clinicians, we finely categorize the sentence-level hallucination outputs of LVLMs into five distinct levels:
@@
Attribute Hallucinations: These manifest as incorrect judgments or descriptions of the size, shape, location, and number of organs and pathologies, and affect diagnostic accuracy to some extent.
Prompt-induced Hallucinations: These hallucinations are induced by prompts containing confused information, often arising from a lack of plausibility or factuality in the prompt, and testing the model’s robustness in specific contexts.
@@ 3.4  Construction of the Benchmark
For images from the VQA dataset, we designed a series of questions based on six dimensions: modality, plane, shape, size, organ, location, and pathology, following the methods proposed in [3]. These original questions are considered conventional medical questions $Q_{Conv}$. We then used the (SOTA) model, LLaVA-Med-pretrained [11], to infer answers to $Q_{Conv}$. The generated answers $A_{Conv}$ were manually evaluated and modified. Correct answers $A_{Conv}^{true}$ were used as ground truth (after correcting imperfect expressions manually), while incorrect answers $A_{Conv}^{false}$ were treated as hallucinatory outputs and their ground truths were re-annotated to expedite the annotation process. Questions of incorrect answers $Q_{Conv}^{false}$ were annotated with the hallucination categories introduced in subsection 3.3.
To quickly expand the dataset, we utilized the GPT-3.5 API to rewrite $Q_{Conv}$ questions, ensuring the questions’ form did not alter the ground truth, resulting in $Q_{Conv}^{true^{\prime}}$ and $Q_{Conv}^{false^{\prime}}$.
@@
In the IRG scenario, we sampled 1800 images and their corresponding medical reports from the MIMIC-test and OpenI datasets (sampling methods are detailed in the Appendix). To generate concise, de-identified medical reports containing complete key information as the dataset’s ground truth, we processed the medical reports by removing sentences comparing the patient’s previous medical history to ensure the accuracy of independent data. We also de-identified the reports by removing patient-specific, doctor-specific, and visit-specific information. After cleaning the data, we concatenated the Findings and Impressions sections of the reports to form the ground truth for the medical image report generation task.
The annotation team consisted of three experienced doctors who conducted evaluations, with a lead doctor responsible for resolving any disagreements. Each question underwent at least two to three rounds of evaluation by each doctor to ensure the quality of the GT. In addition to manually constructing questions based on open-source images, the primary task of the annotators was to review, refine, and correct the model-generated answers used as auxiliary annotations to establish the correct GT. The annotators also labeled the correctness and hallucination type of the model-generated responses
@@ 4  MediHall Score
In traditional NLP tasks, metrics often fail to directly reflect the factual correctness of a language model’s output, typically measuring only shallow similarities to the ground truth. Accuracy, while indicating whether generated content is correct, evaluates at a coarse semantic level and cannot distinguish between degrees of hallucinations in the output. Existing methods like CHAIR and POPE, designed for general LVLM hallucination evaluation, are limited to ’object hallucinations’ in general domains and cannot accommodate the multi-layered complexities of hallucinations in the medical field. Furthermore, they are often constrained to fixed benchmarks or specific types of questions.
@@ 5  MediHallDetector
Figure 2: Visualization of MediHalldetector-related information. (a) Model structure, SFT process, and inference objective of MediHalldetector. (b) Comparison of three rounds of evaluation agreement and average inference time for different evaluation models. (c) Comparison of different evaluation models’ agreement with human evaluation preferences in different hallucination texts.
@@
Framework of MediHallDetector: Figure 2(a) demonstrates the model structure, SFT process and inference objective of MediHallDetector. MediHallDetector is built upon the fundamental architecture of the LLaVA model, incorporating a dual-layer fully connected network with GELU activation functions as the connector. The initial weights for MediHallDetector were taken from LLaVA1.5-7B[16] to leverage its robust foundational capabilities.
Data Preparation: As shown in Figure 2(a), the training data for MediHallDetector consists of three main categories: traditional medical image-text task data, Med-HallMark data, and specific hallucination evaluation instruction pair data. Traditional medical image-text task data, sourced from SLAKE, VQA-RAD, MIMIC-Test and OpenI datasets shown in Figure 1, helps adapt the general LVLM to the medical domain. Med-HallMark data improves the model’s robustness and prevents the assessment model from being induced to produce hallucinations by the hallucination output itself due to the problem of not having seen the source domain. The hallucination evaluation data, manually annotated medical text described in Section 3, helps the model understand different types of hallucinations and follow the instructions.
@@ 6  Experiment and Discussion
model Medical VQA tasks
BertScore METEOR ROUGE-1 ROUGE-2 ROUGE-L BLEU
@@
Accuracy
BLIP2 47.97 16.15 18.98 6.03 17.13 3.46 0.52 0.27
InstructBLIP 35.99 7.47 6.08 0.59 5.3 1.03 0.57 0.34
InstructBLIP13b 36.02 7.59 6.13 0.58 5.32 1 0.58 0.34
LLaVA1.5-7b 54.89 28.33 23.52 9.3 21.16 4.6 0.53 0.28
LLaVA1.5-13b 52.82 25.98 21.52 8.2 19.38 4.18 0.51 0.23
LLaVA-Med (SF) 36.67 8.8 91.17 1.4 9.1 0.03 0.59 0.35
LLaVA-Med (RF) 35.25 6.91 6.34 1.49 6.09 0.54 0.57 0.32
LLaVA-Med (PF) 33.32 3.27 2.85 0.58 2.68 0.06 0.59 0.35
mPLUG-Owl2 55.11 29.39 22.25 8.38 19.77 3.43 0.50 0.24
XrayGPT 44.4 14 10.66 1.17 9.89 0.37 0.36 0.02
mini-gpt4 42.93 12.93 11.14 1.6 10.19 0.39 0.38 0.09
RadFM 43.84 11.81 11.31 2.16 10.68 1.55 0.56 0.32
Table 2: Comparison results of the different models on the IRG tasks in Med-HallMark by various evaluation metrics.
model Medical IRG tasks
BertScore METEOR ROUGE-1 ROUGE-2 ROUGE-L BLEU
@@
BLIP2 33.05 4.88 10.17 1.49 7.66 0.15
InstructBLIP 47.49 13.98 17.56 2.31 13.6 0.73 0.84
InstructBLIP13b 47.47 13.93 17.54 2.34 13.61 0.74 0.84
LLaVA1.5-7b 47.93 11.24 18.77 2.78 14.64 0.72 0.78
LLaVA1.5-13b 47.96 11.8 18.35 2.4 14.49 0.67 0.81
LLaVA-Med (SF) 30.49 0.61 0.27 0 0.27 0.14
LLaVA-Med (RF) 33.28 2.33 2.73 0.32 2.45 0.02
LLaVA-Med (PF) 37.67 1.25 2.24 0.05 2.06 0.09
mPLUG-Owl2 64.49 40.11 32 13.84 28.5 6.79 0.81
XrayGPT 62.62 25.96 27.94 6.59 22.15 3.26 0.79
mini-gpt4 46.43 10.27 15.37 1.75 12.63 0.53 0.88
RadFM 41.18 3.41 5.88 0.76 4.79 0.05
@@ 6.1  Baseline Results on the Med-HallMark
On the IRG task, mPLUG-Owl2 outperforms the other baseline models on most conventional metrics. LLaVA-Meds yields the worst result. The intuitive examples are reflected in the ROUGE-1/2/L with scores close to 0%. A plausible explanation is that LLaVA-Med is full-parameters fine-tuned on the Q&A dataset with short answers, and its inability to perform the IRG task, which requires the giving of long contextual responses. BertScore is considered to be in favourable agreement with human judgment compared to other metrics. In this case, the scores of 64.49% and 62.62% for mPLUG-Owl2 and XrayGPT, respectively. In contrast, BLIP and LLaVA1.5 families achieve comparable performance due to the BertScore of around 47%.
On the MediHall Score metric for assessing hallucinations, we perform corresponding analyses based on different tasks. LLaVA-Med series achieves higher results on the Med-VQA task, which arrive at 0.59, 0.57, and 0.59, respectively, which is essentially on par with the performance of InstructBLIP as well as RadFM, which is 0.58 and 0.56, respectively.
On the IRG task, the LLaVA-Med series, BLIP2, and RadFM cannot produce a computable MediHall Score since the generation format of these models is not suitable for reporting generation scenarios with contextual reasoning properties.
Why IRG scenarios get higher scores: In coarse-grained IRG tasks, models—especially those not specifically trained on IRG data—might generate content that is less clinically relevant. For example, when tasked with describing a chest X-ray, the intent is to have the model provide a comprehensive analysis of the organs, pathologies, and overall health status. However, the model might produce a response such as: "This is a chest X-ray showing a rib cage, lungs, and a heart. The rib cage, lungs, and heart are clearly visible." While this output consists of basic descriptions, from a clinical standpoint, it contains no hallucinations. We cannot classify it as a hallucination simply because it did not provide more detailed information, as that would not align with the definition of a hallucination. Therefore, in such scenarios, if the less significant information aligns with the fundamental facts of the image, it is deemed correct.
Table 3: Comparison of different models in Med-HallMark on Conventional questions $Q_{Conv}^{true^{\prime}}$ ($\clubsuit$’) and Confidence-weakening questions $Q_{Inconfi}$ ($\spadesuit$’).
@@ 6.2  Comparison between different Metrics
Table 4 quantitatively illustrates the strengths and weaknesses of different metrics when evaluating LVLM responses. In various question types, LVLM responses may be correct or exhibit hallucinations. As shown in Figure 3, regardless of whether the LVLM’s answer $A$ aligns with the ground truth $GT$, the METEOR metric shown in Table 4 fails to directly reflect this alignment.
Figure 3: Examples of questions, LVLM answers and GTGT for different types of tasks.
Table 4: Comparison between traditional Metrics and MediHall Score. Each metric corresponds to QA pairs with the same ID in Figure 3
@@
ROUGE score evaluates the presence of matching n-grams or subsequences between two texts. ROUGE-1/2/L matches single words, bigrams, and the longest common subsequence, respectively. Although this can detect some content alignment between $A$ and $GT$, it is prone to extreme cases. For example, in the confidence-weakening question shown in Figure 3(b), the model correctly identifies the largest organ in the image as the lung. However, the $A$ "Lung." fails to match "lung" in $GT$ due to the punctuation, resulting in a failed direct match. Additionally, because $A$ contains only one word, ROUGE-2/L cannot be computed. BLEU has similar issues; without any shared n-grams or subsequences between $A$ and GT, BLEU also scores zero. While BLEU does account for significant length differences between A and GT, making it somewhat more versatile than ROUGE, it still weakly measures factual correctness.
6.3  Evaluations of the MediHallDetector’s Performance